Consider tossing a die and a “success” is getting a 3 or a 6. By writing out the binomial distribution formula, express the probability that in 500 tosses of a die, you see 200 successes. You need not evaluate the expression.
Answer #1

Answer:

Given,

p = 2/6

= 1/3

q = 1 - 1/3

= 2/3

sample n = 500

r = 200

Binomial distribution P(X = r) = nCr*p^r*q^(n-r)

Probability = 500C200*(1/3)^200*(2/3)^300